Andres started biking to the internet cafe traveling 10 mph, after some time the bike got a flat so Andres walked the rest of the way, traveling 6 mph. If the total trip to the internet cafe took 5 hours and it was 38 miles away, how long did Andres travel at each speed?
? hours at 10 mph
? hours at 6 mph

Answer by Boreal(15235)

x hours walking 5-x hours riding
so 6x+10(5-x)=38
6x+50-10x=38
-4x=-12
x=3 hours walking (18 miles) at 6 mph
5-x=2 hours riding (20 miles) at 10 mph